Dry ice is carbon dioxide in the solid state. 0.85 moles of dry ice are placed into a 4.47 L chamber that is maintained at 35 degrees C. What is the pressure in the chamber in atm after the dry ice turns into a gas? Round to two decimal places.
1
Expert's answer
2019-03-20T16:50:27-0400

P(СO2)×V(СO2) = n(СO2) × R × T(СO2)

R = 0.082 L×atm×mol-1×K-1

TK(СO2) = tC(СO2) + 273.15 = (35 + 273.15) K = 308.15 K.

P(СO2) = n(СO2) × R × T(СO2) / V(СO2) = 0.85 mol × 0.082 L×atm×mol-1×K-1 × 308.15 K / 4.47 L = 4.8049 atm ≈ 4.80 atm.